# Designing Antiferromagnetic Spin-1/2 Chains in Janus Fullerene Nanoribbons

### v1

@@ Abstract
We computationally design antiferromagnetic spin-1/2 chains in fullerene nanoribbons by introducing extra C60 cages at one of their edges. The resulting odd number of intermolecular bonds induces an unpaired $\pi$-electron and hence a quantised magnetic moment in otherwise non-magnetic nanoribbons. We further reveal the formation of an antiferromagnetic ground state upon the linear arrangement of spin-$1/2$ C60 cages that is insensitive to the specific structural motifs. Compared with graphene nanoribbons, Janus fullerene nanoribbons may offer an experimentally more accessible route to magnetic edge states with atomic precision in low-dimensional carbon nanostructures, possibly serving as a versatile nanoarchitecture for scalable spin-based devices and the exploration of many-body quantum phases.
@@ I Introduction
II Methodology
Density functional theory (DFT) calculations Hohenberg and Kohn (1964); Kohn and Sham (1965) are performed under the spin-polarised, generalised-gradient approximation (GGA) of Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996), as implemented in the SIESTA package Soler et al. (2002); Artacho et al. (2008); García et al. (2020). A double-$\zeta$ plus polarisation (DZP) basis set is used with an energy cutoff of 400 Ry and a reciprocal space sampling of 10 $k$-points along the periodic direction for structural relaxation. A vacuum spacing in the non-periodic directions larger than 20 Å is adopted throughout all the calculations. Both the lattice constant and atomic positions are fully relaxed using the conjugate gradient method Payne et al. (1992) with a tolerance on forces of 0.02 eV/Å. Our computational model of fullerene nanoribbons contain 660–1020 carbon atoms in the unit cell. For band structure calculations, the $k$-points are increased to 100 to sample the high-symmetry line.
@@ II Methodology
III Results and Discussion
III.1 Crystal structure
III.2 Magnetic moments
For C60 cages in the green circles of Fig. 1(a), however, there are two intermolecular [2 + 2] cycloaddition bonds and three intermolecular C–C single bonds, leaving one unpaired electron. The unpaired electron in each magnetic C60 cage with an odd number of interfullerene bonds has fully quantised spin-1/2, which is confined within the fullerene unit but delocalised at six main carbon sites due to the resonance structure. Fig. 1(c) shows the Schlegel diagram of the six magnetic carbon atoms in their resonance structure, confirming the quasi-localised nature of the unpaired electron (the yellow point). The Mulliken population difference between spin up and down ($\rho_{\uparrow}-\rho_{\downarrow}$) shows that these six carbon atoms contribute to more than 86.8% of the total magnetic moment within the buckyball, i.e., 1 $\mu_{\rm B}$ per magnetic C60.
III.3 Magnetic order
III.4 Electronic structure
III.5 Exchange interactions
III.6 Magnon spectrum
III.7 Robust magnetic edge states
III.8 Experimental feasibility
III.9 Outlook
Our findings envision Janus fullerene nanoribbons as a versatile platform to design quantum spin chains for a wide range of applications from qubit entanglement to spintronics. Because of the intrinsically weak spin–orbit coupling and negligible hyperfine interaction in carbon Droth and Burkard (2013); Slota et al. (2018), the spin coherence time is expected to be very long, making the proposed Janus fullerene nanoribbons promising building blocks for scalable qubit systems. Furthermore, individual spin-1/2 C60 cages can serve as localised quantum dots with controllable exchange interactions. In addition, quantum chains realised by Janus fullerene nanoribbons can enable scalable architectures for quantum simulators, spintronic devices for exploiting spin-polarised currents, and platforms for exploring emergent topological phases and Majorana modes when coupled with superconductors. Overall, the combination of quantised spin, structural flexibility, and long spin coherence places Janus fullerene nanoribbons at the forefront of candidates for next-generation carbon-based quantum technologies.
IV Conclusions
V Acknowledgements

### v2

@@ Abstract
We design antiferromagnetic spin-1/2 chains in fullerene nanoribbons by introducing extra C60 cages at one of their edges. The resulting odd number of intermolecular bonds induces an unpaired $\pi$-electron and hence a quantised magnetic moment in otherwise non-magnetic nanoribbons. We further reveal the formation of an antiferromagnetic ground state upon the linear arrangement of spin-$1/2$ C60 cages that is insensitive to the specific structural motifs. Compared with graphene nanoribbons, Janus fullerene nanoribbons may offer an experimentally more accessible route to magnetic edge states with atomic precision in low-dimensional carbon nanostructures, possibly serving as a versatile nanoarchitecture for scalable spin-based devices and the exploration of many-body quantum phases.
@@ I Introduction
II Results and Discussion
II.1 Crystal structure
II.2 Magnetic moments
For C60 cages in the green circles of Fig. 1(a), however, there are two intermolecular [2 + 2] cycloaddition bonds and three intermolecular C–C single bonds, leaving one unpaired electron. The unpaired electron in each magnetic C60 cage with an odd number of interfullerene bonds has fully quantised spin-1/2, which is confined within the fullerene unit but delocalised at six main carbon sites due to the resonance structure. Fig. 1(c) shows the Schlegel diagram of the six magnetic carbon atoms in their resonance structure, confirming the quasi-localised nature of the unpaired electron (the yellow point). The Mulliken population difference between spin up and down ($\rho_{\uparrow}-\rho_{\downarrow}$) shows that these six carbon atoms contribute to more than 86.8% of the total magnetic moment within the buckyball, i.e., 1 $\mu_{\textrm{B}}$ per magnetic C60.
II.3 Magnetic order
II.4 Electronic structure
II.5 Exchange interactions
II.6 Magnon spectrum
II.7 Robust magnetic edge states
II.8 Experimental feasibility
II.9 Outlook
Our findings envision Janus fullerene nanoribbons as a versatile platform to design quantum spin chains for a wide range of applications from qubit entanglement to spintronics. Because of the intrinsically weak spin–orbit coupling and negligible hyperfine interaction in carbon Droth and Burkard (2013); Slota et al. (2018), the spin coherence time is expected to be very long, making the proposed Janus fullerene nanoribbons promising building blocks for scalable qubit systems. Furthermore, individual spin-1/2 C60 cages can serve as localised quantum dots with controllable exchange interactions. In addition, quantum chains realised by Janus fullerene nanoribbons can enable scalable architectures for quantum simulators, spintronic devices for exploiting spin-polarised currents, and platforms for exploring emergent topological phases and Majorana modes when coupled with superconductors. The same building blocks can also be extended beyond quasi-1D nanoribbons for realising even more complex condensed matter systems in 2D Wu et al. (2025a) such as ferromagnetic Haldane models Pingen et al. (2025) and altermagnetic Shastry-Sutherland lattice Wu et al. (2025b). Overall, the combination of quantised spin, structural flexibility, and long spin coherence places Janus fullerene nanoribbons at the forefront of candidates for next-generation carbon-based quantum technologies.
III Conclusions
Methods
Density functional theory (DFT) calculations Hohenberg and Kohn (1964); Kohn and Sham (1965) are performed under the spin-polarised, generalised-gradient approximation (GGA) of Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996), as implemented in the SIESTA package Soler et al. (2002); Artacho et al. (2008); García et al. (2020). A double-$\zeta$ plus polarisation (DZP) basis set is used with an energy cutoff of 400 Ry and a reciprocal space sampling of 10 $k$-points along the periodic direction for structural relaxation. A vacuum spacing in the non-periodic directions larger than 20 Å is adopted throughout all the calculations. Both the lattice constant and atomic positions are fully relaxed using the conjugate gradient method Payne et al. (1992) with a tolerance on forces of 0.02 eV/Å. Our computational model of fullerene nanoribbons contain 660–1020 carbon atoms in the unit cell. For band structure calculations, the $k$-points are increased to 100 to sample the high-symmetry line. For monolayer qHP networks, the inclusion of the Grimme’s D3 dispersion corrections Grimme et al. (2010) leads to a decrease in lattice constants by merely 0.3% Peng (2022); Shearsby et al. (2025). We therefore neglect the van der Waals interactions hereafter. For electronic structures, we use 100 $k$-points to sample the high-symmetry line.
@@ Methods
Acknowledgements